Express as a single logarithm:
4*lb(M) + (ln(e^6))/2
note:
lb = log base 2
The answer is lb(8M^4), but I don't know how to get there.
Thank you!

Answer by stanbon(75887)

Express as a single logarithm:
4*lb(M) + (ln(e^6))/2
note:
lb = log base 2
--------------
4log2(M) = (1/2)ln(e^6)
---
log2(M^2) = (1/2)*6
---
log2(M^2) = 3
---
M^2 = 2^3
M^2 = 8
M = 2sqrt(2)
====================
Ceers,
Stan H.